Trend Analysis. The following data pertain to PepsiCo, Inc.:

(in millions) Year 2 Year 1

Revenue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . $39,474 $35,137

Net income . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5,658 5,642

Accounts receivable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4,389 3,725

Inventory. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2,290 1,926

Total current assets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10,151 9,130

Total assets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34,628 29,930

Total current liabilities. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7,753 6,860

Total long-term liabilities. . . . . . . . . . . . . . . . . . . . . . . . . . . 9,641 7,702

Total stockholders’ equity . . . . . . . . . . . . . . . . . . . . . . . . . 20,000 18,000

For each item, calculate both the amount in dollars and the percentage change in

Year 2 using trend analysis and Year 1 as the base year.
Answer #1

Trend analysis

Year 2 Year 1 Change in Dollar Change in percentage
Revenue 39474 35137 4337 4337/35137 = 12.34%
Net income 5658 5642 16 16/5642 = 0.28%
Accounts receivable 4389 3725 664 664/3725 = 17.83%
Inventory 2290 1926 364 364/1926 = 18.90%
Total Current assets 10151 9130 1021 1021/9130 = 11.18%
Total assets 34628 29930 4698 4698/29930 = 15.70%
Total Current liabilities 7753 6860 893 893/6860 = 13.02%
Total long term liabilities 9641 7702 1939 1939/7702 = 25.18%
Total Stockholder's equity 20000 18000 2000 2000/18000 = 11.11%